ACCT. 510 – Accounting for Business Decisions
The basic financial statements of a company include the balance sheet, income statement and cash flow statement. Each statement contains specific information about the company. When viewed together, much can be learned about how the company operates and the impact of decisions made during the period.

Obtain the 2017 and 2018 annual reports for Starbucks. Use the information in the financial statements to respond to the following questions:

1. Write out the basic accounting (balance sheet) equation and provide the values reported by Starbucks at September 30, 2018.

2. Did total assets increase or decrease in 2018 over the previous year?

3. How much inventory did the company have at September 30, 2018?

4. What are the major operating revenue accounts? What percentage is each to total revenues?

5. What are the major operating expense accounts? What percentage is each to total operating expenses?

6. Describe the company’s revenue recognition policy, if reported. (See significant accounting policies).

7. What is Starbuck’s gross margin in 2018? Was this an increase or decrease from 2017?

8. What type of costs do you think is included in Starbuck’s ‘Cost of Sales?’

9. How do ‘cost of sales’ differ from ‘operating expenses’

10. What inventory valuation method does the company use?

11. What were cash flows from operating activities?

12. What were cash flows from investing activities?

13. What were cash flows from financing activities?

14. What was the increase/decrease in cash in 2018 from 2017?

15. Ratio Analysis: Compute the following ratios for 2017 and 2018 for Starbucks?

a. Current Ratio

b. Quick Ratio

c. Accounts Receivable Turnover Ratio

d. Inventory Turnover Ratio

e. Debt-to-Equity Ratio

f. Return on Equity

g. Return on Assets

h. Quality of Net Income (Cash Flows from Operating Activities/Net Income)

i. Profit Margin

16. What does each ratio measure and what does your results suggest about the company?

17. If available, find the industry ratios for the most recent year, compare to your results, and discuss why you believe your company differs or is similar to the industry.

18. Were there any economy, industry or company-specific factors that impact Starbuck’s future income and growth? If so, briefly discuss.
Answer #1

1. Balance sheet equation is given by= Total Assets = Total Liabilities+ Total Equity

Values reported by Star bucks as at 30th September 2019 is

Total assets= $ 24,156.4 M

Total Liabilities= 22,980.6 M

Total Equity= 1175.8 M

Balance sheet equation is

$24,156.4 M= 22,980.6M+1,175.8M

2. Did total assets increase or decrease in 2018 over the previous year?

Total assets in 2018 were $24,156.4 M and in 2017 it was 14,365.6 M.

So, Total assets increased by $9,790.8 M

3. How much inventory did the company have at September 30, 2018?

Company have $1,400.5 M inventory at September 30, 2018.

4. What are the major operating revenue accounts? What percentage is each to total revenues?

Company Major revenue accounts are I. Revenue from Company- Operated stores ii. Revenue from Licensed stores

Total revenue= $ 24,719.5

Revenue from Company- Operated stores = $ 19,690.3 M

=19690.3/24719.5= 79.7%

Revenue from Licensed stores= $ 2,652.2 M

Percentage to total revenue= =2652.2/24719.5= 10.7%